OpenStudy (anonymous):

I'm having trouble solving -10t^5+15t^4+9t^3 I got t^3(-10t^2+15t+9) is this correct? not sure how to factor this

OpenStudy (amistre64):

pull out the negative as well

OpenStudy (amistre64):

-t^3( 10t^2-15t-9) now if the quadratic factors nicely there are a few methods to approach it

OpenStudy (anonymous):

so it should be t^3(10t^2+15t+9) ??

OpenStudy (amistre64):

what i do is assume itll factor and set it up as such (t- /10)(t+ /10) then i work out the details 10*9 = 90 we would need to find that factors of 90 that subtract to get 15

OpenStudy (amistre64):

1,90..89 2,45..43 3,30..27 6,15..11 5,18..13 10,9.. 1 unless i missed some factors, this doesnt break apart nicely

OpenStudy (amistre64):

... that and 15-6 = 9, not 11 :) still not 15 tho

OpenStudy (anonymous):

i thought it was set up as (5t-3)(-2t+3) to start i didn't the factor the t's yet.Am i heading in the right direction?

OpenStudy (amistre64):

that will only get your outside terms; not your inside term

OpenStudy (amistre64):

15+6 = 21 , not 15

OpenStudy (amistre64):

we could try the quadratic formula to see what the roots are

OpenStudy (amistre64):

\[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

as I undertand -1 is the gcf? 15=3*5 9=3*3 -10=2*-5 or 5*-2

OpenStudy (amistre64):

\[x=\frac{15\pm\sqrt{15^2-4(10)(-9)}}{2(10)}\] \[x=\frac{15\pm\sqrt{225+360}}{20}\] \[x=\frac{15\pm\sqrt{585}}{20}\] which means its not going to factor into nice little integers for us

OpenStudy (amistre64):

im not sure what a gcf has to do with it ..... but then i am not up to speed with all the methods that are available either

OpenStudy (anonymous):

How do read your previos post?

OpenStudy (amistre64):

its a combonation of 2 values; the +- is regarded as shorthand. when 2 values have the same structure except for the operation; its simpler to combine them with the +-

OpenStudy (amistre64):

otherwise im not sure what you mean by "read"

OpenStudy (anonymous):

without the fraction bar

OpenStudy (amistre64):

15 plus or minus the square root of 585, all divided by 20

OpenStudy (anonymous):

oh thank you for help, the -10 first is throwing off

OpenStudy (anonymous):

throwing me off

OpenStudy (amistre64):

which is why i factored it out with the t^3

OpenStudy (amistre64):

the negative tha tis :)

OpenStudy (anonymous):

after you factor it out with -t^3 then you multiply -t^3 by all coefficients inside the ( ) ?

OpenStudy (amistre64):

if you wanted to put it all back together again , yes. but normally we factor to find roots or zeros or something that is easier to determine in the factored form

OpenStudy (anonymous):

okay thank you for your help :)

OpenStudy (amistre64):

yw